### v2

# Protecting Quantum Information via Many-Body Dynamical Localization

Abstract
Dynamically localized states in quantum many-body systems are fundamentally important in understanding quantum thermalization and have applications in quantum information processing. Here we explore many-body dynamical localization (MBDL) without disorders in a non-integrable quantum XY spin chain under periodical and quadratic kicks. We obtain the localization phase diagram with the MBDL and delocalization states and show dynamical observables to extract the phase diagram. For proper kick strengths in the MBDL regime, we reveal a local dynamical decoupling effect for persistent Rabi oscillation of certain spins. Furthermore, we propose the MBDL-protected quantum information at high temperatures, and present an analysis of the dynamical decoupling to obtain the required system parameters for quantum storage. Compared to other non-thermalized states, the disorder-free MBDL states require much fewer repetitions and resources, providing a promising way to protect and store quantum information robust against thermal noises.
Introduction.— Coherent and controllable many-body quantum dynamics plays an important role in quantum matter engineering and quantum information processing. As a consequence of thermalization, however, coherence in generic many-body quantum systems is fragile once they are excited to a finite energy density above ground states Rigol et al. (2008); D’Alessio et al. (2016). There are several ways to avoid thermalization, apart from integrability. One is the many-body localized (MBL) states induced by sufficiently strong disorders Anderson (1958); Nandkishore and Huse (2015); Abanin et al. (2019); Sierant et al. , and the other is quantum many-body scars Bernien et al. (2017); Turner et al. (2018); Bluvstein et al. (2021); Serbyn et al. (2021); Zhang et al. (2022a), which correspond to the strong and weak breaking of ergodicity, respectively. In MBL systems, all excitations are localized and thermalization is strongly suppressed, which can protect coherence of any initial states. The scar states also have a long coherence time as their dynamics are effectively restricted in a subspace. These localized states may provide resources for quantum information processing Huse et al. (2013); Serbyn et al. (2013a, 2014); Bahri et al. (2015); Yao et al. ; Zhang et al. (2018); Dong et al. (a); Naik et al. (2019); Mi and et al. (2022); Schmid et al. (2024); Yarloo et al. (2024).
Floquet engineering provides another versatile toolbox to understand thermalization and non-equilibrium phases in periodically driven quantum systems D’Alessio and Rigol (2014); Haldar and Das (2017); Moessner and Sondhi (2017); Geier et al. (2021); Shimasaki et al. (2024); Mi and et. al. (2021); Zhang et al. (2022b); Dong et al. (b); Tiwari et al. (2024). As a paradigm of quantum chaos, the quantum kicked rotor (QKR) can avoid chaotic dynamics due to the dynamical localization Fishman et al. (1982); Tian et al. (2004, 2011); Weidemann et al. (2021); Santhanam et al. (2022), which is viewed as Anderson localization in momentum space and has been observed in ultracold atomic gases Moore et al. (1995); Lemarié et al. (2009, 2010); Manai et al. (2015). Increasing efforts are made to study dynamically localized states under many-body interactions, such as interacting QKRs at the few-body limit Adachi et al. (1988); Boness et al. (2006); Rozenbaum and Galitski (2017); Qin et al. (2017); Notarnicola et al. (2018); Chicireanu and Rançon (2021); Paul et al. , integrable case Keser et al. (2016), and mean-field level Shepelyansky (1993); Gligoric et al. (2011); Lellouch et al. (2020), which were experimentally observed using Bose-Einstein condensates in periodically pulsed optical lattices See Toh et al. (2022); Cao et al. (2022). Recent theoretical works suggest that the many-body dynamical localization (MBDL) persists in periodically kicked Bose gases in strongly interacting regimes Rylands et al. (2020); Vuatelet and Rançon (2021, 2023); Fava et al. (2020). The MBDL has just been observed in an interaction-tuable Bose gas Guo et al. . Similar to the dynamical localization, the dynamical freezing of the magnetization for generic initial states in periodically driven spin chains due to emergent conservation laws is predicted Das (2010); Haldar et al. (2018, 2021). Quantum spin chains with tunable disorders and driving fields have also been proposed to realize MBL and scar (confined) states Lazarides et al. (2015); Ponte et al. (2015); Bordia et al. (2017); Sierant et al. (2023); Kormos et al. (2016); Liu et al. (2023); Verdel et al. (2023). The periodically kicked Ising chains with exactly solvable dynamics can serve as a minimal model to study many-body quantum chaos Prosen (2002); Bertini et al. (2018, 2019); Khasseh et al. (2023); Sharma and Bhosale (2024). Nevertheless, the MBDL in clean driven spin systems, which are amenable in quantum processors, and its related coherence protection remain unexplored.
In this work, we investigate the disorder-free MBDL in a non-integrable kicked XY spin chain and its application in protecting quantum information. For the quantum XY chain with a longitudinal field under periodic transverse-field kicks, we obtain the phase diagram with the MBDL and delocalization phases and show dynamical observables to extract the localization phase diagram. We uncover a local dynamical decoupling (DD) effect under proper kick strengths in the MBDL regime, which enables persistent Rabi oscillation of certain spins. Moreover, we demonstrate the MBDL-protected quantum information at high temperatures, and present an analysis of the local DD to obtain the system parameters for robust quantum storage. Our proposal is realizable with superconducting and Rydberg atom quantum simulators.
Model.— We consider a kicked quantum spin chain of length $L$ described by the Hamiltonian:
$H(t)=H_{\text{XY}}+KT\sum_{j=1}^{L}\left(j-j_{0}\right)^{2}\sigma_{j}^{z}\sum_{n=1}^{N_{T}}\delta(t-nT),$ (1)
where $H_{\text{XY}}=\sum_{j=1}^{L}J\left(\sigma_{j}^{+}\sigma_{j+1}^{-}+\sigma_{j+1}^{+}\sigma_{j}^{-}\right)+\Omega\sigma_{j}^{x}$ is the static Hamiltonian with XY interaction between nearest-neighbour spins and a longitudinal field, $K$ and $T$ denote the strength and period of the transverse-field kicks with a quadratic potential centered at $j_{0}$, respectively. Here $\sigma_{j}^{\pm}=\sigma_{j}^{x}\pm i\sigma_{j}^{y}$ with $\sigma_{j}^{x,y,z}$ being the Pauli matrices of $j$-th spin, $J$ is the interaction strength, and $\Omega$ is the strength of the longitudinal field, which leads to coherent spin flips and breaks $U(1)$ symmetry. Notably, we will emphasize in protection of quantum coherence raising from $\Omega$ via the MBDL in this kicked XY chain, which was not considered in kicked Ising chains Prosen (2002); Bertini et al. (2018, 2019); Khasseh et al. (2023); Sharma and Bhosale (2024). Hereafter we consider the high-frequency driving by setting $T=1/16$ and $\Omega=\hbar=1$ as the energy (frequency) unit.
For the kicked spin chain, the Floquet operator $U_{F}$ is
$U_{F}=e^{-iKT\sum_{j=1}^{L}\left(j-j_{0}\right)^{2}\sigma_{j}^{z}}e^{-iH_{\text{XY}}T}\equiv e^{-iH_{F}T},$ (2)
where $H_{F}=\frac{i}{T}\ln U_{F}$ is the effective Floquet Hamiltonian. The wave function at time $t=nT$ after $n$ kick cycles is |$\ket{\psi(t)}=U_{F}^{n}\ket{\psi(0)}$. As the model is non-integrable and $U(1)$-symmetry broken, we study the static and dynamical properties using the exact diagonalization Weinberg and Bukov (2019) for system size $L\leq 14$ and density matrix renormalization group White (1992); Schollwöck (2011); Fishman et al. (2022) for larger $L$. Without loss of generality, we consider centrosymmetric spin chains by setting $j_{0}=(L+1)/2$ under open boundary condition.
Localization phase diagram.— We first study localization properties of many-body Floquet eigenstates {|$\{\ket{\psi^{F}_{\alpha}}\}$. They can be numerically obtained from the eigen equation $U_{F}\ket{\psi^{F}_{\alpha}}=e^{-i\epsilon_{\alpha}T}\ket{\psi^{F}_{\alpha}}$, where $\epsilon_{\alpha}\in[0,2\pi]$ are the folded quasi-energies with $\alpha=\{1,2,..,D\}$ and $D=2^{L}$ as the Hilbert-space dimension. To probe localization, we compute the average inverse participation ratio (IPR) $I_{F}=\frac{1}{D}\sum_{\alpha=1}^{D}\sum_{p=1}^{D}|\braket{p}{\psi_{\alpha}^{F}}|^{4}$, where $|p\rangle$ denote product states in the computational basis. The result of $I_{F}$ in Fig. 1(a) shows the localization phase diagram of Floquet eigenstates in the $K$-$J$ parameter space. For weak (no) kick and strong interaction, the Floquet system is in the delocalization phase with vanishing $I_{F}$ when $L\rightarrow\infty$, which corresponds to thermalization and loss of quantum information of initial states. As $K$ is increased, it can inter the MBDL (dynamical localization) phase with finite $I_{F}$ in the interacting (noninteracting) case. In the localization regime, a proper kick strength can induce the DD of certain spins with persistent Rabi oscillation, such as two edge spins for $L=11$ and $K/\Omega=16\pi/25\thickapprox 2.01$ in Fig. 1(a), which will be discussed later. Here the local DD leads to the decrease of $I_{F}$ (and $O_{F}$) with respect to $K$, as depicted in Fig. 1(c).
Figure 1: (Color online) Localization phase diagrams on the $K$-$J$ parameter space obtained by (a) the IPR $I_{F}$ and (b) the staggered magnetization $O_{F}$ averaging over all Floquet eigenstates of $L=11$. The dashed line in (a) corresponding to the values of $I_{F}\approx 0.009$ is plotted to guide the eye and as a rough estimation of the boundary between the delocalization and MBDL phase regions. A proper kick gives rise to the DD of certain spins. The dashed lines with triangles and circles for critical points in (b) represent the phase boundaries determined by finite-size scaling of $O_{F}$ and half-chain entanglement $S_{F}$ of Floquet eigenstates, respectively. (c) $I_{F}$ and $O_{F}$ as a function of $K$ near the DD. (d) $S_{F}$ versus $L$ for various $K$. (e) $\bar{r}$ as a function of $K$ for various $L$. Other parameter is $J=1$ in (c) and $J=2$ in (d,e).
Another quantity to reveal localization in the kicked spin chain is the infinite-time staggered magnetization evolving from an initial Néel state $|\ket{\psi(0)}=|\uparrow\downarrow\uparrow\downarrow...\rangle$, which can be calculated via the Floquet ensemble {|$\{\ket{\psi_{\alpha}^{F}}\}$ as $O_{F}=\frac{1}{L}\sum_{j=1}^{L}\sum_{\alpha=1}^{D}(-1)^{j}|c_{\alpha}|^{2}\bra{\psi^{F}_{\alpha}}\sigma^{z}_{j}\ket{\psi^{F}_{\alpha}}$ Rigol et al. (2008), with the coefficients $c_{\alpha}=\braket{\psi^{F}_{\alpha}}{\psi(0)}$. Figure 1(b) shows the result of $O_{F}$ consistent with the phase diagram in Fig. 1(a). The phase boundary can be estimated by the finite-size scaling of $O_{F}$ (see Fig. 5 in the Supplementary Material (SM) SM ). We also compute the average half-chain entanglement entropy: $S_{F}=-\frac{1}{D}\sum_{\alpha=1}^{D}\operatorname{Tr}\left[\rho^{F}_{\alpha}\log\rho^{F}_{\alpha}\right]$, where $\rho^{F}_{\alpha}=\operatorname{Tr}_{j\leqslant L/2}[\ket{\psi^{F}_{\alpha}}\bra{\psi^{F}_{\alpha}}]$ with even $L$. The finite-size scaling of $S_{F}$ Tang et al. (2021); SM gives roughly consistent boundary between the delocalization and MBDL phases in Fig. 1(b). Figure 1(d) shows the area-law entanglement in the deep MBDL phase since only the area of the bipartitioning contributes to the entanglement Page (1993); Luitz et al. (2015); Kjäll et al. (2014); Pal and Huse (2010), which is different from the volume law in the delocalization phase.
Figure 1(e) indicates the crossover from delocalization to the MBDL by the mean gap-ratio parameter $\bar{r}$ D’Alessio and Rigol (2014); Zhang et al. (2020). The level-spacing statistics of Floquet eigenstates is characterized by $r_{\alpha}=\min\{\delta_{\alpha},\delta_{\alpha+1}\}/\max\{\delta_{\alpha},\delta_{\alpha+1}\}$, where $\delta_{\alpha}=\epsilon_{\alpha+1}-\epsilon_{\alpha}$ is the quasi-energy gap. We obtain $\bar{r}$ up to $L=14$ by averaging $r_{\alpha}$ over all Floquet eigenstates and over 20 realizations of spin chains with a center offset $j_{0}\rightarrow j_{0}+j_{\text{offset}}$ (to lift the level-spacing degeneracy), with $j_{\text{offset}}$ uniformly distributed in the small region $[-0.02,0.02]$. The MBDL phase is indicated by a Poissonian level-spacing distribution with $\bar{r}\approx 0.386$, whereas presents a Wigner-Dyson distribution with $\bar{r}\approx 0.53$ in the delocalized phase. It remains an important task to capture the delocalization-MBDL transition and criticality in the thermodynamics limit Rispoli et al. (2019). However, our results shown in Fig. 1 pinpoint the large parameter region for the MBDL phase in the kicked spin chain.
Dynamical observables.— The (de)localization can be observed via the stroboscopic dynamics of the kicked spin chain. For an initial product state in typical experiments, a direct observation is the spin-up (excitation) density distribution $P_{\uparrow}(j,t)=\left\langle\psi(t)\left|\frac{\sigma^{z}_{j}+\mathbb{I}}{2}\right|\psi(t)\right\rangle$ at the stroboscopic time $t=nT$, where $\mathbb{I}$ is the $2\times 2$ identity matrix. The results of $P_{\uparrow}(j,t)$ for different initial states show similar dynamics SM : The spin chain tends to thermalize in the static case. Turning on periodical kicks up to the MBDL regime, the quantum coherence of initial states preserves without thermalization in the long-time evolution. This disorder-free non-thermalized dynamical phase is due to the kinetic constriction by the quadratic kick potential. The MBDL can be broken down and the thermalization dynamics dominates as $J$ is increased.
Figure 2: (Color online) (a) Time evolution of $\mathcal{I}$ for typical parameters. (b) $\overline{\mathcal{I}}$ as a function of $J$ for $L=13,15,19$. (c) Phase diagram extracted by $\overline{\mathcal{I}}$. (d) Time evolution of subsystem entanglement $S_{A=L/2}$ for various parameters. (e) Saturation entanglement $\bar{S}$ versus $L$. Other parameters are $L=13$ in (a,e) and $K=0.5$ in (b,d,e), and the initial Néel state is $|\ket{\psi(0)}=|\uparrow\downarrow\uparrow\downarrow...\rangle$ in (a-e).
To better characterize the MBDL and extract the localization phase diagram, one can prepare an initial Néel state and measure the evolution of the generalized imbalance $\mathcal{I}(t)=\frac{1}{L}\sum_{j}(-1)^{j+1}P_{\uparrow}(j,t)$. Typical results of $\mathcal{I}(t)$ are shown in Fig. 2(a). In the MBDL (delocalization) phase when $L\rightarrow\infty$, $\mathcal{I}$ remains finite (becomes vanishing) after long time evolution, which is expected to tend to the value of $O_{F}$ in the infinite-time limit. We can use the long-time-averaged imbalance $\overline{\mathcal{I}}$ (averaging over the time interval from $t=100T$ to $200T$ in simulations) to probe (de)localization and extract the phase diagram, as shown in Figs. 2(b) and 2(c), respectively.
Another observable is the dynamics of subsystem entanglement entropy $S_{A}(t)=-\operatorname{Tr}\left[\rho_{A}(t)\log\rho_{A}(t)\right]$ Kaufman et al. (2016); Rispoli et al. (2019); Dong et al. (a); Choi et al. (2023), where $\rho_{A}(t)=\operatorname{Tr}_{j\leq A}[\ket{\psi(t)}\bra{\psi(t)}]$ is the reduced density matrix of subsystem-A. By performing tomography, one can measure $\rho_{A}$ after $n$ kicks to obtain $S_{A}(t)$. Figure 2(d) shows distinct saturation behaviors of the half-chain entanglement $S_{A=L/2}$ for various $L$ in the MBDL and delocalization phases. In finite localized (delocalized) systems, $S_{A=L/2}(t)$ approaches to saturate at late times to certain value independent (linearly dependent) of $L$ Fagotti and Calabrese (2008); Bardarson et al. (2012); Serbyn et al. (2013b); Hart et al. (2021); Lim et al. (2024). The system-size dependence of the saturation entanglement $\bar{S}$, numerically obtained by averaging $S_{A=L/2}(t)$ over the time interval from $t=400T$ to $600T$, is shown in Fig. 2(e). The result confirms the area- and volume-law of the half-chain entanglement in the MBDL and delocalization phases, respectively.
MBDL-protected quantum information.— In the deep MBDL phase, arbitrary states in the energy spectrum are non-thermalized. Due to this strong breaking of ergodicity, the quantum information can be robustly encoded with initial MBDL states. For a generic initial state $|\psi(0)\rangle$, its effective inverse temperature $\beta_{\text{eff}}$ can be obtained by solving $\operatorname{Tr}[H_{F}(\ket{\psi(0)}\bra{\psi(0)}-\rho_{\beta})]=0$, where $\rho_{\beta}=e^{-\beta_{\text{eff}}H_{F}}/\operatorname{Tr}(e^{-\beta_{\text{eff}}H_{F}})$ is the thermal density matrix Bañuls et al. (2011); Chen et al. (2021). We plot $\beta_{\text{eff}}$ as a function of quasi-energies of all product states $\epsilon_{p}=\bra{\psi(0)}H_{F}\ket{\psi(0)}$ for the system of $L=11$ in the MBDL phase and local DD case in Fig. 3(a) and 3(b), respectively. To show the protection of quantum information at high temperatures, we can consider the initial product states with $\beta_{\text{eff}}\approx 0$ near the center of the energy spectra. The non-thermalization dynamics of initial MBDL states can be observed from the long-time coherent evolution of spin density distributions, with an example shown in Fig. 3(d) (and the SM SM ). We can study additional information on the coherent dynamics beyond local observables. We compute the time evolution of fidelity $F_{A}(t)=\left[\operatorname{Tr}\sqrt{\sqrt{\rho_{A}(t)}\rho_{A}(0)\sqrt{\rho_{A}(t)}}\right]^{2}$ and entanglement entropy $S_{A}(t)$ of subsystem-A with $A=4$ for the high-temperature initial states in Figs. 3(e) and 3(f). The subsystem fidelity with high values and low entanglement entropy persists in the long-time evolution.
Figure 3: (Color online) $\beta_{\text{eff}}$ as a function of $\epsilon_{p}$ for (a) MBDL with $K=1$; and (b) DD with $K=16\pi/25$. The initial high-temperature states with $\beta_{\text{eff}}\approx 0$ are denoted in red. (c) Time evolution of $F_{A}$ and $S_{A}$ (with $A=4$) for the initial state in (a). Time evolution of (d) $P_{\uparrow}$; and (e) $F_{A}$, $S_{A}$ and $\langle\sigma_{1}^{z}\rangle$ for the initial state in (b). (f) Lifetime $\tau$ of the $\langle\sigma_{1}^{z}\rangle$-oscillation versus $L$ for the initial vacuum and entangled states SM . Other parameters are $J=0.5$ and $L=11$.
[…]
Figure 4: (Color online) (a) $\omega_{d}$ with respect to $\Omega$ for the DD of edge spins in the chain of $L=11$. (b) $K_{d}^{(1,1)}$ with respect to $L$. (c) $K$ versus $N_{d}$ for $L=21,41,61$. (d) Evolution of $P_{\uparrow}(j,t)$ in the DD case of $N_{d}=4$ for $K=16\pi/25$ and $L=21$.
Analysis of local DD.— We proceed to present an analysis of the local DD to obtain the required system parameters. For the $j$-th decoupled spin in the MBDL regime, its effective coupling strength is vanishing, corresponding to the local Floquet operator $U^{(j)}_{F}=e^{-i\tilde{K}\sigma_{j}^{z}}e^{-i\Omega T\sigma_{j}^{x}}$, where $\tilde{K}=KT(j-j_{0})^{2}$. When $\tilde{K}=m\pi$ with $m$ being positive integers for certain kick strengths, the operator reduces to $U^{(j)}_{F}=(-1)^{m}e^{-i\Omega T\sigma_{j}^{x}}$. This implies the Rabi oscillation of the effectively decoupled spin with the frequency $\omega_{d}=\Omega T/\pi$. The analytical expression of the oscillation frequency is consistent of the numerical result, as shown in Fig. 4(a). The kick strength required for the DD of $j$-th spin is $K_{d}^{(j,m)}=m\pi/T(j-j_{0})^{2}$. For the centrosymmetric chain of odd spins considered here, the local DD occurs with pairs of spins, such as $j=1,L$ for $L=11$ shown in Fig. 3(d) and Fig. 4. We numerically extract $K_{d}^{(1,1)}$ for various $L$ from the local dynamics and find well agreement with the analytical expression $K_{d}^{(1,1)}=4\pi/T(L-1)^{2}$, as shown in Fig. 4(b).
To store more quantum information without thermalization, one needs more decoupled spins. Remarkably, we obtain the following relation between the kick strength and the DD spin number $N_{d}$ of system size $L$ SM :
$K(N_{d},L)=\frac{\pi}{T\lfloor(L-1)/N_{d}\rfloor^{2}},$ (3)
[…]
Discussions and conclusions.— Before concluding, we discuss experimental realization of the kicked XY quantum spin chain. The first system is the superconducting quantum simulators, consisting of 1D arrays of controllable qubits with tunable nearest-neighbour XY coupling Li et al. ; Dong et al. (b); Tan et al. (2021). One can individually and dynamically tune local potentials to realize the static and periodical kick terms along the $x$- and $z$-axis in Eq. (1), similar as those in experiments in Ref. Dong et al. (b). Alternatively, one can realize the Floquet unitary $U_{F}$ given by Eq. (2) in a digital way using tunable quantum gates on superconducting qubits Mi and et. al. (2021); Zhang et al. (2022b). The second platform is the 1D driven Rydberg atom array Bluvstein et al. (2021); Chen et al. (2023). One can encode the effective spin-1/2 in a pair of opposite-parity Rydberg states and realize the XY spin Hamiltonian via the resonant dipole-dipole interactions Chen et al. (2023). The kicked potential may be produced by applying pulsed laser beams with spatially dependent light shifts Bluvstein et al. (2021); Chen et al. (2023). Both superconducting and Rydberg quantum simulators are programmable with controllable many-body dynamics to detect the MBDL phase and the related coherence protection.
In summary, we have explored the disorder-free MBDL in the non-integrable quantum XY chain under periodical kicks. We have revealed the localization phase diagram with dynamical observables and large parameter regions for the MBDL, and uncovered the local DD effect for persistent Rabi oscillation of certain spins. We have proposed the MBDL-protected quantum information at high temperatures, and presented an analysis of the DD to obtain the required system parameters for quantum storage. The MBDL states in spin systems offer a promising platform for studying quantum thermalization and quantum information processing robust against thermal noises.
Acknowledgements.
The authors acknowledge stimulating discussions with Chushun Tian, Dong-Ling Deng, Zhen-Yu Wang, and Shi-Liang Zhu. This work was supported by the National Natural Science Foundation of China (Grants No. 12174126 and No. 92365206), Innovation Program for Quantum Science and Technology (No. 2021ZD0301802), the Guangdong Basic and Applied Basic Research Foundation (Grant No. 2024B1515020018), the Science and Technology Program of Guangzhou (Grant No. 2024A04J3004), and Quantum Science Center of Guangdong-Hong Kong-Macau Great Bay Area.